# Send Message to the Future? Blockchain-based Time Machines for Decentralized Reveal of Locked Information

### v3

@@ Abstract
Conditional information reveal systems automate the release of information upon meeting specific predefined conditions, such as time or location. This paper introduces a breakthrough in the understanding, design, and application of conditional information reveal systems that are highly secure and decentralized. By designing a new practical timed-release cryptography system and a secret sharing scheme with reveal-verifiability, a novel data sharing system is devised on the blockchain that ‘sends messages in the future’ with highly accurate decryption times. Notably, the proposed secret sharing scheme applies to other applications requiring verifiability of revealed secret shares. This paper provides a complete evaluation portfolio of this pioneering paradigm, including analytical results, a validation of its robustness in the Tamarin Prover and a performance evaluation of a real-world, open-source system prototype deployed across the globe. Using real-world election data, we also demonstrate the applicability of this innovative system in e-voting, illustrating its capacity to secure and ensure fair electronic voting processes.
@@ I Introduction
This paper introduces a smart-contract-based timed-release cryptography system with improvements in security and efficiency compared to the existing proposals [8, 9, 10, 11, 12]. Moreover, this paper shows how general conditional information reveal systems can be secured using the same architecture as the proposed timed-release cryptography system.
@@
Notably, this paper also proposes a secret sharing scheme with reveal-verifiability for conditional information reveal systems to enhance their efficiency and security. The common definition of verifiable secret sharing (VSS) and publicly verifiable secret sharing (PVSS) in existing literature [13] only provides verifiability of the generation and dissemination of secret shares. However, in timed-release cryptography, verifiability of the revealed secret shares is also required to ensure the reconstructed secret is the one sent by the client. This novel secret sharing scheme introduces this property named reveal-verifiability, that provides public verifiability in the secret reconstruction stage.
@@
  • A smart-contract-based architecture to achieve secure conditional information reveal in a decentralized way. A secure timed-release cryptography system is constructed under this architecture.
@@
The rest of the paper is structured as follows: Section 2 reviews the related work in conditional information reveal systems and timed-release cryptography. Section 3 introduces a secure architecture for conditional information reveal systems using smart contracts and secret sharing. Section 4 presents a smart-contract-based timed-release cryptography system. Section 5 presents a new cryptographic protocol that secures the proposed system. Section 6 shows the advantages of the proposed system compared to the existing methods. Section 7 showcases the applicability of the proposed system in an e-voting scenario and Section 8 concludes the paper with an outline of future work.
@@ II-A2 The Secret Holders Approach
Centralized designs [1, 33, 34, 35, 36, 37, 38] rely on one independent party acting as a secret holder. These methods focus on cryptographic-level security under the assumption of the honesty and activeness of the single secret holder. Security and robustness at a system level are not considered, such as potential single point of failure, and high level of trust required for the secret holder. The decentralized methods [39, 2, 40, 41, 42] utilize threshold cryptography to provide fault tolerance while involving multiple secret holders to guard a secret.
@@
The development of blockchains and smart contracts addresses the challenges of incentives, data availability and authenticity. Recent proposals [8, 9, 10, 11, 12] use smart contracts that integrate cryptocurrencies to provide incentives for secret holders, and ensure secret shares are immutable and available after revealing. However, none of these proposals provides undeniable reveal-verifiability, that is, secret shares revealed by honest secret holders can be claimed as incorrect shares in the presence of malicious clients. The possibility of having malicious clients that frame secret holders is only mentioned by Ning et al. [9], but a solution to prevent false claims of incorrect shares is not proposed. Other proposals [8, 10, 11, 12] do not consider malicious clients in their security models.
@@
Decentralized on-demand (smart-contract-based): [8, 9, 10, 11, 12] Yes Yes No Yes No
@@
The decentralized on-demand approach only provides services upon receiving clients’ requests. The existing proposal [2] is closest to our approach as our approach also falls into this category. However, it does not offer incentives for honesty, resistance to time source poisoning, or Sybil attack resistance. Incentives and Sybil attack resistance are resolved in the smart-contract-based proposals [8, 9, 10, 11, 12], but there lack of discussion on more secure time references of decryption time. More importantly, they also fall short in terms of reveal-verifiability under malicious clients, that is, malicious clients can construct some incorrect reveal-verification to frame an honest secret holder revealing incorrect secret share.
@@ II-B Secret Sharing Schemes for Secret Holders Approach
The original Shamir’s Secret Sharing scheme does not provide verifiability to the correctness of secret shares when sharing to the secret holders [43]; proposals on verifiable secret sharing [44, 45] provides verifiability in the distribution phase to ensure that holders receive correct shares.
However, it is not noticed in the existing work on the secret holder approach that verifiability in the distribution phase is not enough for timed-release cryptography. The message recovery stage where secret holders publish their secret shares is essential for timed-release cryptography systems, otherwise, it introduces attack opportunities for adversarial holders to submit incorrect shares to fail message recovery or for adversarial clients to make malicious claims that holders submit incorrect shares. To the best of our knowledge, secret sharing schemes with reveal-verifiability have not been demonstrated.
@@ II-B1 Required Background
III Towards Secure Conditional Information Reveal Systems
Taking a step back from timed-release cryptography, this section introduces a paradigm to show that smart contract platforms can be applied to serve generic conditional information reveal systems with key security advantages.
III-A Security Challenges
  1. 1.
  2. 2.
  3. 3.
  4. 4.
III-B Decentralizing Conditional Information Reveal Systems
Considering the smart-contract-based proposals in time-based conditional information reveal system [8, 9, 10, 11, 12] in a generalized view, this architecture can in fact be applied to all types of conditional information reveal systems. The only difference is that the secret holders require different oracles to check if the required conditions are met.
In general, involving multiple secret holders through secret sharing and communicating through smart contracts improves system reliability. To elaborate, the advantages of such an architecture include:
Data authenticity and availability
Information holders and clients interact with smart contracts deployed on a blockchain network. This form of communication offers key advantages such as immutability and non-repudiation of messages, which are crucial for maintaining accountability and ensuring honest behavior among all parties involved in the system.
Fault tolerance and Sybil attack resistances
Involving multi-secret-holders with secret sharing protocols provides fault tolerance and robustness for the system, but it raises the risks of Sybil attacks in the permissionless setting where anyone can become a secret holder. Incorporating smart contracts for cryptocurrency staking addresses this issue.
Network autonomy through encoded rules in the smart contract
The smart contract provides a platform for secret holders or the wider public to raise disputes transparently. Network autonomy is achieved by disputing secret holders and using protocols for verifiable information sharing. For example, secret holders can dispute the actions of other holders via the smart contract. Sufficient disputes against a secret holder results in blacklisting and loss of cryptocurrencies as a punishment.
Figure 2: Modules in the decentralized architecture
IV A Smart-Contract-based Timed-Release Cryptography System
Going back into time-based conditional information reveal, this section illustrates a timed-release cryptography system using the decentralized conditional information reveal architecture. Sharing a similar basic idea with other smart-contract-based constructions [8, 9, 10, 11, 12], Our system additionally achieves higher reliability and security by making uses of the blockchain clock and achieving reveal-verifiability to prevent malicious clients from framing honest secret holders.
IV-A The Smart-Contract-based Architecture
Figure 3: The high-level process of sending a timed-release message
IV-B Improved Clock Reliability with On-Chain Timestamps
IV-C Verifiability of Revealed Information
To ensure the messages are reconstructed correctly by the secret holders and prevent malicious behavior from both holders and clients, a novel cryptographic protocol is proposed. This protocol addresses a limitation found in existing on-demand timed-release cryptography schemes based on secret holders, where the correctness of the shared information is not publicly verifiable.
In the proposed system, the new cryptographic protocol introduces a modified version of Shamir’s Secret Sharing to provide public verifiability for the secret shares submitted by the holders. Similar to existing secret holder systems that utilize Shamir’s Secret Sharing, each holder is assigned a unique secret share. The reveal of the secret is possible if the number of secret holders that publish their shares reaches a threshold.
The key advancement of the proposed protocol lies in providing verifiability to ensure the correctness of secret shares throughout the dissemination and reconstruction stage effectively with low communication costs. This also ensures fair reward distribution to honest holders based on the correctness of their submitted secret shares. The details of this novel cryptographic protocol are elaborated in the subsequent section.
IV-D Example Workflow
(a) 4 honest secret holders
(b) 3 honest holders and 1 adversarial holder
Figure 4: Example workflow
Overall, the proposed system ensures that the secret remains confidential until the specified decryption time. It also incentivizes honest behavior from the holders through the deposit and reward mechanisms.
V Secret Sharing with Reveal-Verifiability on Public Communication Channel
@@
V-A Objectives and Problem Modelling
The objective of the cryptographic protocol is to hide, distribute, and later recover a piece of information $k$. When applied to the timed-release cryptography system, $k$ is a symmetric key to encrypt and decrypt a time-sensitive message $m$. Two roles are involved in this protocol, on the one hand, it is the client that holds $k$, on the other hand, multiple secret holders receive shares of $k$ for recovery in the future.
Unlike existing secret sharing protocols, since the proposed architecture is blockchain-based, it is required that clients and secret holders only communicate via sending permissionless blockchain transactions, i.e. in a publicly available broadcasting channel. Moreover, verifiability of revealed secret shares is an additional requirement compared to existing secret sharing methods.
We assume all honest parties are not compromised, i.e., secret holders keep their secret key safe, and clients keep their secret information $k$ safe. The broadcasting channel is assumed to be reliable such that all messages are eventually delivered to all parties. The order of the messages can be changed because each secret share is independent and can be revealed in arbitrary order.
V-B Design
@@
Figure 5: A secret sharing example with four holders
@@
V-C Applying to Timed Release Cryptography
@@ VI Security Analysis
In this section, an attack model for the proposed system is outlined, followed by providing formal proof of security within this framework. Moreover, a comparative analysis of different secret holder approaches is provided to demonstrate the advantage of the proposed system on critical security features.
@@ VI-A Attack Model
As the foundation of the proposed system, the integrity and data availability of the underlying blockchain are assumed to hold, thereby guaranteeing the safe transmission of data from clients and secret holders. Additionally, the randomly generated values including private keys $sk$ and the clients’ random values $r$ are assumed to be secure and confidential.
Three types of adversaries are taken into account in the attack model, external adversaries, malicious secret holders, and malicious clients. External adversaries are malicious parties that aim to obtain confidential information using public knowledge, specifically, the encrypted messages before decryption time.
In a distributed and permissionless setting, the existence of malicious secret holders is considered, while assuming a majority of holders remain benign. This assumption can be upheld through various measures, tailored to specific application contexts. For instance, in a voting scenario, voters themselves could act as secret holders, while in more general applications, monetary incentives via cryptocurrency could encourage honest behavior. Under these conditions, whether a minority of malicious secret holders could compromise the secrecy of a time lock message is assessed.
Malicious clients also have the potential to disrupt the system if they can compromise the verifiability of decrypted messages. The system fails if clients can falsely accuse honest secret holders of submitting incorrect shares. In this case, honest secret holders are punished and disqualified even if they follow the protocol. Therefore, the proposed protocol must ensure verifiability of secret shares in the existence of malicious clients.
VI-B Message Secrecy Before Decryption
Three lemmas are provided to prove message secrecy before $t$ secret shares are published. Lemma VI.1 and VI.2 combined is a proof by induction showing that message secrecy holds against external adversaries regardless of the number of secret holders in the system. Lemma VI.3 shows that message secrecy also holds against a minority of malicious secret holders.
@@ VI-B Message Secrecy Before Decryption
Lemma VI.1.
For all generated and secure parameters of holders’ secret keys $sk$ and the clients’ choice of $r$, an external adversary can never use the public knowledge $(g_{1},g_{2},pk_{1},pk_{2},pk_{3},g_{1}^{r},g_{2}^{r},\alpha_{3})$ to derive a secret share before the corresponding holders submit it.
@@
The lemma is successfully proven by the Tramarin Prover. Source code can be found in the open-source repository [14] of this project. ∎
@@
Lemma VI.2.
An external adversary cannot obtain any secret share $s_{i}$ before its respective holder submits it, irrespective of the total number of secret holders.
@@
For an adversary to access a secret share $s_{i}$ before holder $i$ submits it in a system with $n>3$ holders ($i<n$), it must obtain a secret share prematurely from $n-1$ holders, as the additional knowledge provided by the $n^{th}$ holder, $pk_{n}=g_{1}^{sk_{n}}$ and $\alpha_{n}=g_{1}^{sk_{n}*r}\oplus P(n)$, does not provide any information about $s_{i}=g_{1}^{sk_{i}*r}$. Similarly, the adversary must also be able to prematurely access a secret share with $n-2$ holders. Therefore, the adversary must always be able to access a secret share without the presence of one holder. When the number of secret holders is decreased to $n=3$, the adversary cannot obtain a secret share as proven in Lemma VI.1. Hence, an external adversary cannot prematurely access a secret share before the holder submits it at any system size of $n$. ∎
@@
Lemma VI.3.
For a system with any number of secret holders $n\geq 3$ and threshold of $t>n/2$, $t-1$ malicious secret holders cannot obtain $k$ before an honest secret holder submits its share.
@@
Suppose the collaborating malicious secret holders can obtain $k$ before honest holders submit shares, they must be able to gain one extra secret share belonging to the honest holders. That is, they are able to get the secret share using the public knowledge $g_{1},g_{2},g_{1}^{r},g_{2}^{r},pk_{1},...,pk_{n},\alpha_{t+1},...,\alpha_{n}$ plus their own secret keys. Since their own secret keys are randomly generated values, which are irrelevant to any other honest holders’ secret share, they must be able to obtain the secret share using public knowledge. However, lemma VI.2 shows that it is not possible to obtain a secret share using public knowledge, which contradicts the initial assumption and proves the lemma. ∎
VI-B1 Verifiability of Correct Decryption
In the verification process, an honest secret holder should always pass the verification even if the encryption request is initiated by a malicious client.
Based on the process of shares distribution and verification, each client publishes a value $a=g_{1}^{r}$ for secret holders to compute the share as $s=a^{sk}$ and another value $b=g_{2}^{r}$ for the public to verify secret shares. The verification process is reduced to checking whether $e(pk,b)=e(s,g_{2})$ holds. However, malicious client may not follow the protocol on the computation of $a$ and $b$.
Lemma VI.4.
For a malicious client publishing any $a$ and $b$, and a secret holder computing and submitting $s=a^{sk}$, any other party can always identify that $s=a^{sk}$ holds without knowing $sk$, or identify that the client is dishonest.

### v4

@@ Abstract
Conditional information reveal systems automate the release of information upon meeting specific predefined conditions, such as a designated time in the future. By designing a new practical timed-release cryptography architecture and a secret sharing scheme with verifiable information reveal, a novel data sharing system is devised on smart contracts that “sends messages in the future” with highly accurate decryption times. Using the same architecture, this paper also introduces a breakthrough in the understanding, design, and application of conditional information reveal systems that are highly secure and decentralized. A complete evaluation portfolio is provided to this pioneering paradigm, including analytical results, a validation of its robustness in the Tamarin Prover and a performance evaluation of a real-world, open-source system prototype deployed across the globe. Using real-world election data, we also demonstrate the applicability of this innovative system in e-voting, illustrating its capacity to secure and ensure fair electronic voting processes.
@@ I Introduction
This paper introduces a smart-contract-based timed-release cryptography system with improvements in security and efficiency compared to the existing proposals [8, 9, 10, 11, 12, 13]. Moreover, this paper shows how general conditional information reveal systems can be secured using the same architecture as the proposed timed-release cryptography system.
@@
Notably, this paper proposes a secret sharing scheme with reveal-verifiability for conditional information reveal systems to enhance their efficiency and security. The common definition of verifiable secret sharing (VSS) and publicly verifiable secret sharing (PVSS) in existing literature [14] only provides verifiability of the generation and dissemination of secret shares. However, in timed-release cryptography, verifiability of the revealed secret shares is also required to ensure the reconstructed secret is correct. This novel secret sharing scheme introduces such a property named reveal-verifiability, that provides public verifiability in the secret reconstruction stage.
@@
  • A secure timed-release cryptography system is constructed using smart contract and the proposed secret sharing scheme. The same architecture is applicable to achieve secure conditional information reveal systems in general.
@@
The rest of the paper is structured as follows: Section 2 reviews the related work in conditional information reveal systems and timed-release cryptography. Section 4 presents a smart-contract-based timed-release cryptography system. Section 3 presents a new secret sharing protocol that secures the proposed system. Section 4 presents a smart-contract-based timed-release cryptography system. Section 5 applies the proposed architecture to secure general conditional information reveal systems using smart contracts and secret sharing. Section 6 shows the advantages of the proposed system compared to the existing methods. Section 7 showcases the applicability of the proposed system in an e-voting scenario and Section 8 concludes the paper with an outline of future work.
@@ II-A2 The Secret Holders Approach
Centralized designs [1, 34, 35, 36, 37, 38, 39] rely on one independent party acting as a secret holder. These methods focus on cryptographic-level security under the assumption of the honesty and activeness of the single secret holder. Security and robustness at a system level are not considered, such as potential single point of failure, and high level of trust required for the secret holder. The decentralized methods [40, 2, 41, 42, 43, 8, 9, 10, 11, 12, 13] utilize threshold cryptography to provide fault tolerance while involving multiple secret holders to guard a secret.
@@
The development of blockchains and smart contracts addresses the challenges of incentives, data availability and authenticity. Recent proposals [8, 9, 10, 11, 12, 13] use smart contracts that integrate cryptocurrencies to provide incentives for secret holders, and ensure secret shares are immutable and available after revealing. However, none of these proposals provides reveal-verifiability. The definition of reveal-verifiability for secret sharing schemes is presented below.
Definition 1.
Reveal-verifiability: Let $k$ be a secret distributed among $n$ participants using a secret-sharing scheme, and let $R$ represent the reconstruction transcript, which includes the $t$ pieces of information used for reconstructing the secret. The reconstruction process is defined by a deterministic reconstruction function $\mathcal{R}$:
$k^{*}=\mathcal{R}(R),$
where $R$ is the reconstruction transcript comprising the $t$ secret shares used for reconstruction, and $k^{*}$ is the output of the reconstruction function.
Then, a secret-sharing scheme satisfies reveal-verifiability if there exists a deterministic verification function $\mathcal{V}$, accessible to all participants, such that:
$\mathcal{V}(R)=\begin{cases}\text{true},&\text{if }\mathcal{R}(R)=k,\\ \text{false},&\text{otherwise}.\end{cases}$
The reveal-verifiability problem
To ensure a secret share revealed by a secret holder is correct, certain information as the witness is required. For example, Bacis et al. [12] propose that a client publish the cryptographic hashes of the secret shares as witnesses. This allows the revealed shares to be hashed and compared against the witnesses for verification. However, this approach assumes the client is honest and does not account for the possibility of a malicious client publishing false witnesses. If a client publishes random hash values, the revealed secret shares are mistakenly deemed incorrect, leading to punishment of honest secret holders and compromising the integrity of the system. While malicious clients are modeled by Ning et al. [9], but a solution to prevent such cases is not proposed. Other proposals [8, 10, 11, 12, 13] similarly fail to address the threat of malicious clients, leaving them free to publish false witnesses without detection.
@@
Decentralized on-demand (smart-contract-based): [8, 9, 10, 11, 12, 13] Yes Yes No Yes No
@@
The decentralized on-demand approach only provides services upon receiving clients’ requests. The existing proposal [2] is closest to our approach as our approach also falls into this category. However, it does not offer incentives for honesty, resistance to time source poisoning, or Sybil attack resistance. Incentives and Sybil attack resistance are resolved in the smart-contract-based proposals [8, 9, 10, 11, 12, 13], but there lack of discussion on more secure time references of decryption time. More importantly, they also fall short in terms of reveal-verifiability in the presence of malicious clients.
@@ II-B Secret Sharing Schemes for Secret Holders Approach
The original Shamir’s Secret Sharing scheme [44] lacks a mechanism to verify the correctness of the secret shares distributed to the secret holders. While verifiable secret sharing protocols [45, 46] introduce verification in the distribution phase to ensure that holders receive valid shares, they do not address the need for verifiability in the reveal phase, where the correctness of the recovered secret must be confirmed. Recent secret sharing protocols for smart-contract-based timed-release cryptography systems introduce various variants of secret sharing, for example, Li and Palanisamy [11] proposed an interactive protocol that enables secret holders to select points on the polynomial to use as shares without revealing them to the client; Jiang et al. [13] incorporate identity-based encryption into the secret sharing process. However, these variants still fail to achieve reveal-verifiability, allowing malicious clients to frame honest secret holders in the reveal stage.
@@ II-B1 Required Background
III Secret Sharing with Reveal-Verifiability on Public Communication Channel
@@
III-A Problem Modeling and Assumptions
The objective of the cryptographic protocol is to distribute preserve confidentiality of a fixed-length information $k$. When applied to the timed-release cryptography system, $k$ is a symmetric key to encrypt and decrypt a time-sensitive message $m$. Two roles are involved in this protocol, including the client that owns $k$ and the $n$ secret holders that receive secret shares of $k$ denoted as $s_{1},...,s_{n}$. Here we define this required confidentiality property:
Definition 2.
Confidentiality: For any probabilistic polynomial-time (PPT) adversary $\mathcal{A}$, the adversary’s advantage in obtaining $k$ fewer than $t$ shares is negligible. That is:
$\Pr\big{[}\mathcal{A}(s_{1},s_{2},\ldots,s_{t-1})\to k\big{]}\leq\epsilon,$
where $\epsilon$ is a negligible function, $s_{1},s_{2},\ldots,s_{t-1}$ are the shares available to the adversary, and $t$ is the reconstruction threshold of the secret-sharing scheme.
Unlike existing secret sharing protocols, since the proposed architecture is blockchain-based, it is required that clients and secret holders only communicate via sending permissionless blockchain transactions, i.e. in a publicly available broadcasting channel. Moreover, reveal-verifiability (Definition 1) is an additional requirement compared to existing secret sharing methods.
We make the following cryptographic and system level assumptions as the security foundations of the proposed protocol:
  • We assume the discrete logarithm problem is hard, that is, Let $G_{1}$ and $G_{2}$ be cyclic groups of prime order $p$ with generators $g_{1}$ and $g_{2}$, respectively. Given $g_{1}$ and $g_{1}^{x}$, it is computationally infeasible for a probabilistic polynomial-time adversary to compute $x\in\mathbb{Z}_{p}$.
  • At the system level, we assume at least a threshold $t$ out of $n$ secret holders are honest.
  • The randomly generated values, including private keys $sk_{1},...,sk_{n}$ from secret holders and the clients’ random values $r$ (introduced below), are assumed to be secure and confidential.
  • The broadcasting channel is assumed to be reliable such that all messages are eventually delivered to all parties.
III-B Design
@@
Figure 2: An example of the proposed secret sharing method with four secret holders
@@
III-C Applying to Timed Release Cryptography
@@
IV A Smart-Contract-based Timed-Release Cryptography System
This section illustrates a timed-release cryptography system using the decentralized conditional information reveal architecture. Sharing a similar basic idea with other smart-contract-based constructions [8, 9, 10, 11, 12, 13], Our system additionally achieves higher reliability and security by making uses of the blockchain clock and achieving reveal-verifiability to prevent malicious clients from framing honest secret holders.
IV-A The Smart-Contract-based Architecture
Figure 3: The high-level process of sending a timed-release message
IV-B Improved Clock Reliability with On-Chain Timestamps
IV-C Verifiability of Revealed Information
To ensure the messages are reconstructed correctly by the secret holders and prevent malicious behavior from both holders and clients, the proposed secret sharing protocol is adopted.
IV-D Example Workflow
(a) 4 honest secret holders
(b) 3 honest holders and 1 adversarial holder
Figure 4: Example workflow in the view of on-chain activities
V Towards Secure Conditional Information Reveal Systems
Taking a step back from timed-release cryptography, this section introduces a paradigm to show that smart contract platforms and the proposed secret sharing scheme also apply to generic conditional information reveal systems with key security advantages.
V-A Security Challenges
  1. 1.
  2. 2.
  3. 3.
  4. 4.
V-B Decentralizing Conditional Information Reveal Systems
Considering the smart-contract-based proposals in time-based conditional information reveal system [8, 9, 10, 11, 12, 13] in a generalized view, this architecture can in fact be applied to all types of conditional information reveal systems. The only difference is that the secret holders require different methods to check if the required conditions are met.
Involving multiple secret holders through secret sharing and communicating through smart contracts improves system reliability. It provides a reliable communication channel where data authenticity and availability are ensured. Combining incentive mechanisms with cryptocurrencies and the proposed secret sharing protocol that provides reveal-verifiability, honest and active secret holders are always incentivized to secure the protocol.
Figure 5: Modules in the decentralized architecture for conditional information reveal systems using smart contract and secret sharing
@@ VI Security Analysis
Going back into the proposed timed-release cryptography system, this section outlines an attack model for the proposed system and provides formal proof of security within this framework. Moreover, a comparative analysis of different secret holder approaches is provided to demonstrate the advantage of the proposed system on critical security features.
@@ VI-A Attack Model
Three types of adversaries are taken into account.
  • External adversaries are malicious parties that aim to breach confidentiality (see Definition 2) using public knowledge.
  • Malicious secret holders are parties that aim to breach confidentiality using public knowledge along with their secret shares. We assume at most $t-1$ secret holders are considered malicious. This assumption can be upheld through various measures, tailored to specific application contexts. For instance, in a voting scenario, voters themselves could act as secret holders, while in more general applications, monetary incentives via cryptocurrency could encourage honest behavior. Under these conditions, whether a minority of malicious secret holders could compromise the secrecy of a time lock message is assessed.
  • Malicious clients are parties that aim to breach reveal-verifiability (see Definition 1) of the protocol. The system fails if clients can falsely accuse honest secret holders of submitting incorrect shares. In this case, honest secret holders are punished and disqualified even if they follow the protocol. Therefore, the proposed protocol must ensure reveal-verifiability in the presence of malicious clients.
VI-B Confidentiality
Three lemmas are provided to prove the confidentiality of the proposed protocol. Lemma 1 and 2 combined is a proof by induction showing that confidentiality holds against external adversaries regardless of the number of secret holders in the system. Lemma 3 shows that confidentiality also holds against a minority of malicious secret holders.
@@ VI-B Confidentiality
Lemma 1.
When $n=3,t=2$, For any PPT external adversaries $\mathcal{A}$ that have access to the public knowledge $(g_{1},g_{2},pk_{1},pk_{2},pk_{3},g_{1}^{r},g_{2}^{r},\alpha_{3})$ and at most one secret share $(s_{i},i\in\{1,2,3\})$, the probability of knowing $k$ is negligible.
@@
Suppose $\mathcal{A}$ can obtain $k$ with the public knowledge and an $s_{i}$ with non-negligible probability, $\mathcal{A}$ must be able to get an additional secret share $s_{j}=pk_{j}^{r}$. Then $\mathcal{A}$ must be able to get $r$ from $g_{1}^{r}$ or $g_{2}^{r}$ with non-negligible probability, that is, $\mathcal{A}$ must be able to solve the discrete logarithm problem with non-negligible probability, which contradicts to our assumption.
The lemma is also successfully proven by the Tamarin Prover. Source code can be found in the open-source repository [15] of this project. ∎
@@
Lemma 2.
For any $n>3$, any PPT external adversaries $\mathcal{A}$ has negligible probability of obtaining any secret share $s_{i}$ before its respective holder submits it.
@@
For $\mathcal{A}$ to access a secret share $s_{i}$ before holder $i$ submits it in a system with $n>3$ holders ($i<n$), it must obtain a secret share prematurely from $n-1$ holders. The additional knowledge provided by the $n^{th}$ holder is $pk_{n}=g_{1}^{sk_{n}}$ and $\alpha_{n}=g_{1}^{sk_{n}*r}\oplus P(n)$. Based on the hardness of discrete logarithm problem, $\mathcal{A}$ has negligible probabilities to obtain $sk_{n}$ and $r$, and therefore has negligible probability to obtain $s_{i}=g_{1}^{sk_{i}*r}$. Similarly, the adversary must also be able to prematurely access a secret share with $n-2$ holders. Therefore, the adversary must always be able to access a secret share without the presence of one holder. When the number of secret holders is decreased to $n=3$, the adversary cannot obtain a secret share as proven in Lemma 1. Hence, an external adversary has a negligible probability of prematurely accessing a secret share before the holder submits it at any system size of $n$. ∎
@@
Lemma 3.
Any $t-1$ PPT malicious secret holders has negligible probability of knowing $k$.
@@
Suppose the collaborating malicious secret holders can obtain $k$ before honest holders submit shares, they must be able to gain one extra secret share belonging to the honest holders. That is, they are able to get the secret share using the public knowledge $g_{1},g_{2},g_{1}^{r},g_{2}^{r},pk_{1},...,pk_{n},\alpha_{t+1},...,\alpha_{n}$ along with their own secret keys. Since their own secret keys are randomly generated values, which are irrelevant to any other honest holders’ secret share, they must be able to obtain the secret share using public knowledge. However, lemma 2 shows that it is not possible to obtain a secret share using public knowledge, which contradicts the initial assumption and proves the lemma. ∎
VI-C Reveal-Verifiability
Shares published by honest secret holders should be verified as correct even if the encryption request is initiated by a malicious client. Below we present possible actions of a malicious client and show that reveal-verifiability is not violated.
Denote $a=g_{1}^{r}$ for secret holders to compute the share as $s=a^{sk}$ and another value $b=g_{2}^{r}$ for the public to verify secret shares. In this notation, the verification process is to check whether $e(pk,b)=e(s,g_{2})$ holds. A malicious client can attempt to violate the protocol by publishing $a$ or $b$ as any other values other than $a=g_{1}^{r}$ and $b=g_{2}^{r}$.
Lemma 4.
There exists a deterministic verification function $\mathcal{V}$, accessible to all participants, that identifies $a=g_{1}^{r}$ and $b=g_{2}^{r^{\prime}}$ for any $r\neq r^{\prime}$.